# A Cascaded Dilated Convolution Approach for Mpox Lesion Classification

### v2

@@ Abstract
The global outbreak of the Mpox virus, classified as a Public Health Emergency of International Concern (PHEIC) by the World Health Organization, presents significant diagnostic challenges due to its visual similarity to other skin lesion diseases. Traditional diagnostic methods for Mpox, which rely on clinical symptoms and laboratory tests, are slow and labor intensive. Deep learning-based approaches for skin lesion classification offer a promising alternative. However, developing a model that balances efficiency with accuracy is crucial to ensure reliable and timely diagnosis without compromising performance. This study introduces the Cascaded Atrous Group Attention (CAGA) framework to address these challenges, combining the Cascaded Atrous Attention module and the Cascaded Group Attention mechanism. The Cascaded Atrous Attention module utilizes dilated convolutions and cascades the outputs to enhance multi-scale representation. This is integrated into the Cascaded Group Attention mechanism, which reduces redundancy in Multi-Head Self-Attention. By integrating the Cascaded Atrous Group Attention module with EfficientViT-L1 as the backbone architecture, this approach achieves state-of-the-art performance, reaching an accuracy of 98% on the Mpox Close Skin Image (MCSI) dataset while reducing model parameters by 37.5% compared to the original EfficientViT-L1. The model’s robustness is demonstrated through extensive validation on two additional benchmark datasets, where it consistently outperforms existing approaches.
@@ I Introduction
While early applications of deep learning to Mpox detection show promise, they face several critical methodological challenges. One major issue is the considerable variation in model development and validation, with many studies relying on evaluations using only a single dataset. A model that performs well on one dataset may not deliver similar results on others, which limits its generalizability and reduces its reliability when applied to diverse real-world data.
@@
The development of such models must also carefully balance computational efficiency with diagnostic accuracy. While faster processing times are desirable, they cannot come at the cost of reduced accuracy, as reliable diagnosis remains the primary goal in clinical settings.
Despite the demonstrated effectiveness of transformer-based architectures in skin lesion classification [20][21], their application to Mpox detection has been minimal. Most existing studies focus predominantly on traditional Convolutional Neural Networks , leaving a significant gap in leveraging the capabilities of transformer models for this specific task.
@@ II-B Deep Learning for Mpox Classification
Figure 2: Integration of CAGA Module within the Network Architecture (left) and detailed framework of Cascaded Group Attention with Cascaded Atrous Attention Module (right). Left: The MBConv layer represents the Mobile Inverted Bottleneck Convolution[44], which is part of the EfficientViT architecture. A Depthwise Separable Convolution[30] is used to reduce computational complexity while preserving spatial and channel-wise feature interactions, converting the input to an $n\times h$ dimension. Batch Normalization (BN)[45] is applied to the output of the residual connection.
@@
While these prior works have established the viability of deep learning for Mpox detection, they demonstrate a common trade-off between model complexity and performance, often requiring either resource-intensive ensemble approaches or simplified architectures that may compromise accuracy. This study builds upon these foundations by taking a fundamentally different approach and leveraging the strengths of both CNN and transformer architectures. Rather than focusing solely on architectural depth or ensemble methods, I address the core challenge of efficient feature extraction through the novel Cascaded Atrous Attention mechanism, achieving high accuracy across multiple datasets while staying computationally efficient.
@@ III Methodology
This section presents the Cascaded Atrous Group Attention framework. I first provide a comprehensive overview of the framework, followed by detailed explanations of its key components: the Cascaded Atrous Attention (CAA) module and the Cascaded Group Attention (CGA) mechanism.
@@ III-A Cascaded Atrous Group Attention (CAGA)
I introduce Cascaded Atrous Group Attention, (see Fig. 2 (right)) an architecture that builds upon the Cascaded Group Attention (CGA)[46] framework. While the CGA directly processes each attention head using standard self-attention, this work replaces this mechanism with the proposed Cascaded Atrous Attention (CAA). By integrating multi-scale dilated convolutions and a sophisticated self-attention mechanism for each dilation, CAA enhances the feature representation capabilities of the CGA approach.
@@ III-A1 Cascaded Atrous Attention (CAA)
introduce a hierarchical cascading approach to dilated self-attention maps that progressively aggregates and enhances spatial context. The model incrementally integrates proximal and broader neighborhood features, creating a multi-scale representation that captures nuanced spatial dependencies across different contextual granularities. To address dimensionality disparities between the dilated attention maps and subsequent dilated convolutional inputs, I implement a $1\times 1$ convolution projection layer, thereby ensuring dimensional consistency.
@@ V Conclusion
In this work, I proposed the Cascaded Atrous Group Attention framework, which effectively combines the Cascaded Atrous Attention module and the Cascaded Group Attention mechanism to improve the efficiency and accuracy for Mpox classification. CAA achieved multi-scale representation using dilated convolutions and cascaded these outputs to enhance spatial and contextual information. CAA was encapsulated in a CGA, originally designed to address the redundancy in Multi-Head Self-Attention. Experimental results show that the approach highlighted state-of-the-art results on three datasets. In future work, I plan to extend the application of CAGA to broader tasks, such as disease segmentation and object detection, where multi-scale feature representation and efficient attention mechanisms could drive substantial improvements. Integrating self-supervised learning techniques with CAGA could improve its performance, especially when applied to limited or imbalanced medical datasets.

### v3

@@ Abstract
The global outbreak of the Mpox virus, classified as a Public Health Emergency of International Concern (PHEIC) by the World Health Organization, presents significant diagnostic challenges due to its visual similarity to other skin lesion diseases. Traditional diagnostic methods for Mpox, which rely on clinical symptoms and laboratory tests, are slow and labor intensive. Deep learning-based approaches for skin lesion classification offer a promising alternative. However, developing a model that balances efficiency with accuracy is crucial to ensure reliable and timely diagnosis without compromising performance. This study introduces the Cascaded Atrous Group Attention (CAGA) module to address these challenges, combining the Cascaded Atrous Attention and the Cascaded Group Attention mechanisms. The Cascaded Atrous Attention module utilizes dilated convolutions and cascades the outputs to enhance multi-scale representation. This is integrated into the Cascaded Group Attention mechanism, which reduces redundancy in Multi-Head Self-Attention. By integrating the Cascaded Atrous Group Attention module with EfficientViT-L1 as the backbone architecture, this approach achieves state-of-the-art performance, reaching an accuracy of 98% on the Mpox Close Skin Image (MCSI) dataset while reducing model parameters by 37.5% compared to the original EfficientViT-L1. The model’s robustness is demonstrated through extensive validation on two additional benchmark datasets, where it consistently outperforms existing approaches.
@@ I Introduction
While early applications of deep learning to Mpox detection show promise, they face several critical methodological challenges. One major issue is the considerable variation in model validation, with many studies relying on evaluations using only a single dataset. A model that performs well on one dataset may not deliver similar results on others, which limits its generalizability and reduces its reliability when applied to diverse real-world data.
@@
A critical gap exists in Mpox classification models that are both rigorously tested and strike a balance between computational efficiency and diagnostic accuracy. While reducing processing times is important, it should never compromise the accuracy of the diagnosis, as reliable diagnosis remains the primary goal in clinical settings.
@@ II-B Deep Learning for Mpox Classification
Figure 2: Integration of CAGA within the Network Architecture (left) and detailed structure of Cascaded Group Attention with Cascaded Atrous Attention Module (right). Left: The MBConv layer represents the Mobile Inverted Bottleneck Convolution[42], which is part of the EfficientViT architecture. A Depthwise Separable Convolution[28] is used to reduce computational complexity while preserving spatial and channel-wise feature interactions, converting the input to an $n\times h$ dimension. Batch Normalization (BN)[43] is applied to the output of the residual connection.
@@
These works in Mpox detection highlight a trade-off between model complexity and performance, often relying on resource-intensive ensembles or simplified architectures. This study addresses the core challenge of efficient feature extraction with the novel Cascaded Atrous Group Attention mechanism, achieving high accuracy across multiple datasets while remaining computationally efficient.
@@ III Methodology
This section presents the Cascaded Atrous Group Attention. I first provide a comprehensive overview of the module, followed by detailed explanations of its key components: the Cascaded Atrous Attention (CAA) and the Cascaded Group Attention (CGA) mechanisms.
@@ III-A Cascaded Atrous Group Attention (CAGA)
I introduce Cascaded Atrous Group Attention (see Fig. 2 (right)) an architecture that builds upon the Cascaded Group Attention (CGA)[44] module. While the CGA directly processes each attention head using standard self-attention, this work replaces this mechanism with the proposed Cascaded Atrous Attention (CAA). By integrating multi-scale dilated convolutions and a sophisticated self-attention mechanism for each dilation, CAA enhances the feature representation capabilities of the CGA approach.
@@ III-A1 Cascaded Atrous Attention (CAA)
I introduce a hierarchical cascading approach to dilated self-attention maps that progressively aggregates and enhances spatial context. The model incrementally integrates proximal and broader neighborhood features, creating a multi-scale representation that captures nuanced spatial dependencies across different contextual granularities. To address dimensionality disparities between the dilated attention maps and subsequent dilated convolutional inputs, I implement a $1\times 1$ convolution projection layer, thereby ensuring dimensional consistency.
@@ V Conclusion
In this work, I proposed the Cascaded Atrous Group Attention architecture, which combines the Cascaded Atrous Attention and the Cascaded Group Attention mechanism to improve the efficiency and accuracy for Mpox classification. CAA achieved multi-scale representation using dilated convolutions and cascaded these outputs to enhance spatial and contextual information. CAA was encapsulated in a CGA, originally designed to address the redundancy in Multi-Head Self-Attention. Experimental results show that the approach highlighted state-of-the-art results on three datasets. In future work, I plan to extend the application of CAGA to broader tasks, such as disease segmentation and object detection, where multi-scale feature representation and efficient attention mechanisms could drive substantial improvements. Integrating self-supervised learning techniques with CAGA could improve its performance, especially when applied to limited or imbalanced medical datasets.